3.) Calculate the pH of a 0.398M sodium hydroxide solution. (This will take 2 steps). - VII
Answer #1

NaOH,being a strong base will dissociate completely in aqueous solution.

NaOH \rightleftharpoons Na+ + OH-
Here, 1 M of NaOH will give 1 M of Na+ and 1 M of OH-

Similarly, 0.398 M of NaOH will give 0.398 M of Na+ and 0.398 M of OH-

We know,

pOH = - log [OH-]

pOH = - log(0.398)

pOH = 0.400

We also know,

pOH+ pH = 14
pH = 14 - pOH

pH = 14 - 0.400

pH = 13.6